OpenStudy (anonymous):

solve: log4(x+1) + log4(x-1) = 6. I'm not sure how to get the 6 to a log4 base... thanks!

Nnesha (nnesha):

first look at some rules for log\[\huge\rm \log ~x + \log~ y = \log (x \times y)\] \[\large\rm log ~x - \log~ y = \log \frac{ x }{ y }\] so there is plus sign how would you change that ??

OpenStudy (anonymous):

log4[(X+1)(x-1)] = 6 ? it's the 6 that's throwing me off..

Nnesha (nnesha):

haha okay you have to change logarithmic to exponential form

Nnesha (nnesha):

 right side number become exponent of base and move left side number to the right side like tin this example